OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively

OpenStudy (nikato):

So now u know the geometric sequence formula?

OpenStudy (nikato):

Well it's an= a(r) ^ n-1 Where a=1st term R=ratio N= number of terms

OpenStudy (anonymous):

@nikato i get that but how do i do it if i dont have the first term

OpenStudy (nikato):

Ok, u want to solve though system of equations. Did u plug everything it so u have to different equations?

OpenStudy (anonymous):

im honestly so confused

OpenStudy (nikato):

Ok, so just follow along

OpenStudy (anonymous):

Thank you so much~

OpenStudy (nikato):

Plug in the numbers into the geometric formula 768=ar^4 -12=ar^1 Ok with this?

OpenStudy (anonymous):

Where did you get the ^4 from?

OpenStudy (nikato):

So in the geoetric formula it's n-1. Since its the 5th term, it's 5-1 or 4

OpenStudy (anonymous):

oh okay okay

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

So now I'm gonna divide both equations

OpenStudy (nikato):

768=ar^4 ------------ -12=ar^1

OpenStudy (anonymous):

divided them by eachother

OpenStudy (nikato):

Remember a/a=1 and ur just subtracting exponents when u divide So this equals -64=r^3 Can u find r?

OpenStudy (nikato):

And yes, I divided them by each other

OpenStudy (anonymous):

no :/

OpenStudy (nikato):

So u want to cube root both sides. Which u can do on ur calculator

OpenStudy (anonymous):

-4

OpenStudy (nikato):

Or u can factor 64 which is 4x4x4 but becuz it's negative it's -4x-4x-4

OpenStudy (nikato):

Ok. Yes. So our r=-4

OpenStudy (anonymous):

how do we find a?

OpenStudy (nikato):

So now let's plug it back into one of the original equation. Lets use the second one -12=ar^1 So -12=a(-4)^1 -12=a(-4) So now, can u find a?

OpenStudy (anonymous):

i got -3/4

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

3

OpenStudy (anonymous):

would it be an=3*-4^n-1

OpenStudy (nikato):

Yea, correct. Great job!

OpenStudy (anonymous):

Thank you so much i actually get it now!

OpenStudy (nikato):

Ur welcome!